# Incremental Rotation Averaging Revisited and More: A New Rotation Averaging Benchmark

### v3

@@ III-B Other Key Steps Stemming from the IRAv3+
Local absolute rotation global alignment and optimization: Given the alignment rotation of each cluster, the absolute rotations in the cluster’s local coordinate system are globally aligned to the reference’s coordinate system. After that, inlier edge set of the original EG is firstly obtained based on the aligned rotations and then used for providing constraints for globally optimizing them (cf. Eq. 11 and Eq. 12).
IV New Rotation Averaging Benchmark Creation
It could be observed that most of the existing rotation averaging methods [26, 27, 28, 29, 16, 21, 30, 19, 31, 22, 23, 24] perform performance evaluation based on the 1DSfM [32] dataset with the slightly outdated Bundler’s [33] camera calibration results serving as the ground-truth source of camera rotations, which has been presented for more than ten years. In order to give a more comprehensive evaluation on the existing mainstream rotation averaging methods, a new 1DSfM-derived benchmark based on COLMAP [6] is rebuilt here. COLMAP is a general-purpose, end-to-end image-based 3D reconstruction pipeline with a graphical and command-line interface, which is proposed more recently and updated more frequently. In recent years, COLMAP has become the most widely recognized and commonly used open-source SfM toolbox with various downstream applications, such as novel view synthesis [38, 39] and neural surface reconstruction [40, 41]. As a result, we believe that it is time for COLMAP to take the place of Bundler for rotation averaging benchmarking. The new benchmark creation process mainly contains three parts: epipolar-geometry graph regeneration, ground-truth data acquisition, and evaluation metrics definition, which are thoroughly described in the following, respectively.
IV-A Epipolar-Geometry Graph Regeneration
In order to present a new rotation averaging benchmark by making full use of COLMAP, the Epipolar-geometry Graph (EG) of each test data in the 1DSfM dataset is regenerated at first. It should be noted that only the images in the Maximum Connected Component (MCC) are considered during the EG regeneration procedure here for the efficiency consideration, and this image connectivity information is provided by the file contained in the original 1DSfM dataset. Then, the images to be processed are fed into the COLMAP toolbox for local feature extraction and matching with default parameters. Given the pair-wise feature matches produced by COLMAP, OpenCV library is employed for essential matrix estimation and decomposition to produce relative motion (rotation and translation), where the feature-level point to epipolar line distance threshold and the desirable level of estimation confidence are set to and respectively during the RANdom SAmple Consensus (RANSAC) [42] procedure of the essential matrix estimation [10] process. After that, the image pair with more than pairs of inlier feature matches for RANSAC-based essential matrix estimation is added as a graph edge to the regenerated EG. It should be further noted that the camera intrinsic parameters are required during essential matrix estimation, which are provided by the file contained in the original 1DSfM dataset here.
Data | |||||||||||||
ALM | |||||||||||||
ELS | |||||||||||||
GDM | |||||||||||||
MDR | |||||||||||||
MND | |||||||||||||
NYC | |||||||||||||
PDP | |||||||||||||
PIC | |||||||||||||
ROF | |||||||||||||
TOL | |||||||||||||
TFG | |||||||||||||
USQ | |||||||||||||
VNC | |||||||||||||
YKM |
IV-B Ground-Truth Data Acquisition
After EG regeneration, COLMAP is employed again for the ground-truth data acquisition of camera absolute rotations (and locations) for rotation averaging evaluation. Specifically, for each test data of the 1DSfM dataset, given the image matching result of COLMAP itself obtained in the last subsection, the SfM procedure integrated in COLMAP is conducted with default parameters, and the camera calibration result is denoted as , where is the set of cameras successfully registered by COLMAP. In addition, the camera calibration result of Bundler, provided by the file contained in the original 1DSfM dataset, is denoted as . It should be noted that has been aligned to the GPS coordinate system and is with real physical scale, while suffers from the issue of scale ambiguity due to the metric reconstruction nature of SfM. As a result, in order to use the camera calibration result of COLMAP as ground-truth camera poses for rotation (and translation) averaging evaluation, it should be aligned to the coordinate system of Bundler for scale recovery. The alignment is achieved by estimating a similarity transformation [43] based on RANSAC with the co-calibrated cameras by COLMAP and Bundler, , where the locational distance threshold and the desirable estimation confidence are set to and , respectively. Then, based on the estimated transformation , could be aligned to by:
Before performing rotation averaging method benchmarking, it should be further noticed that the basic assumption of involving the camera calibration result of the incremental SfM methods (Bundler or COLMAP) to serve as ground-truth source for motion (rotation and translation) averaging evaluation lies in that the camera orientation and localization results of incremental SfM with iterative RANSAC-based outlier filtering and BA-based parameter optimization [44] are acknowledged to be with much higher accuracy than those of motion averaging before performing the final global BA. Nonetheless, the incremental SfM-based calibration result inevitably contains estimation errors and is still not the real ground-truth camera poses. To make the evaluation with more reliability, inspired by the work of long-term visual localization benchmarking [45], the camera poses calibrated by COLMAP that pass Bundler-based cross check with certain pose accuracy level are used as the final camera pose ground-truth data. Its basic idea lies in that the camera poses calibrated by both COLMAP and Bundler with enough closeness in the -dimensional camera pose space are used for the following rotation averaging evaluation. Specifically, given the rotational () and locational () distance thresholds with different accuracy levels of camera poses: high-accuracy , medium-accuracy , and coarse-accuracy , the camera subsets for rotation averaging evaluation with respect to the different distance thresholds are obtained by:
The sizes of the camera subsets together with some other meta-data about the new rotation averaging benchmark are listed in Table I, and in order to make a trade-off between ground-truth data magnitude and reliability, the camera subset with medium-accuracy, i.e. , are finally employed for rotation averaging benchmarking. The COLMAP-to-Bundler alignment median errors of rotation and location in Table I are defined as:
where returns the median value of the set . It could be further observed from the table that: 1) in most cases, except for MDR, TOL, and YKM, more cameras could be registered by COLMAP than Bundler; 2) the disparity in camera calibration accuracy between COLMAP and Bundler is insignificant as both and are relatively small; 3) Different test data in the new benchmark is with different degree of conformity in accuracy () between COLMAP and Bundler, take for example, co-calibrated cameras pass the median-level-accuracy cross check for MND, while for TFG, that value decreases to only .
Data | IRLS-GM | IRLS- | MPLS | DESC | HARA | NeuRoRA | IRA | IRA++ | IRAv3 | IRAv3+ | IRAv4 | MPLS | IRAv4 |
[35] | [26] | [27] | [31] | [30] | [28] | [21] | [22] | [23] | [24] | w/ BA | w/ BA | ||
ALM | |||||||||||||
ELS | |||||||||||||
GDM | |||||||||||||
MDR | |||||||||||||
MND | |||||||||||||
NYC | |||||||||||||
PDP | |||||||||||||
PIC | |||||||||||||
ROF | |||||||||||||
TOL | |||||||||||||
TFG | |||||||||||||
USQ | |||||||||||||
VNC | |||||||||||||
YKM | |||||||||||||
Rank |
IV-C Evaluation Metrics Definition
For a particular rotation averaging method (what could be is described in the next section), it is firstly performed on the relative rotations from the regenerated EG in this paper to obtain the absolute camera rotations. Then, based on the rotation averaging result and the relative translation measurements, translation averaging is performed with the mainstream method, BATA555https://bbzh.github.io/ [34], to get the absolute camera locations, and the -based motion averaging result is denoted as . Then, the rotation and similarity transformation for aligning to the coordinate system of are estimated by the methods of voting-based single rotation averaging [29] and RANSAC-based similarity estimation, respectively. The pose alignment is conducted in the same way of Eq. 13, and the aligned poses are denoted as . Finally, the median rotational and locational errors on the camera subset of are served as the evaluation metrics of the new rotation averaging benchmark, which are defined as:
V Rotation Averaging Method Benchmarking
In this section, several mainstream rotation averaging methods for comparison are first introduced, and then rotation averaging method benchmarking in terms of rotation and location estimation is performed based on the presented new rotation averaging benchmark.
Data | IRLS-GM | IRLS- | MPLS | DESC | HARA | NeuRoRA | IRA | IRA++ | IRAv3 | IRAv3+ | IRAv4 | MPLS | IRAv4 |
[35] | [26] | [27] | [31] | [30] | [28] | [21] | [22] | [23] | [24] | w/ BA | w/ BA | ||
ALM | |||||||||||||
ELS | |||||||||||||
GDM | |||||||||||||
MDR | |||||||||||||
MND | |||||||||||||
NYC | |||||||||||||
PDP | |||||||||||||
PIC | |||||||||||||
ROF | |||||||||||||
TOL | |||||||||||||
TFG | |||||||||||||
USQ | |||||||||||||
VNC | |||||||||||||
YKM | |||||||||||||
Rank |
V-A Rotation Averaging Methods for Comparison
Except for the IRA series (IRA [21], IRA++ [22], IRAv3 [23], IRAv3+ [24], and IRAv4 proposed in this paper) which have been described in Sec. II and Sec. III, there are some other mainstream rotation averaging methods involved in this section for comparison, including IRLS-GM666https://ee.iisc.ac.in/cvlab/research/rotaveraging/ [35], IRLS-6 [26], MPLS777https://github.com/yunpeng-shi/MPLS [27], DESC888https://github.com/ColeWyeth/DESC [31], HARA999https://github.com/sunghoon031/HARA [30], and NeuRoRA101010https://github.com/pulak09/NeuRoRA [28]. These methods are briefly introduced in the following for better understanding.
IRLS-GM [35] works on the Lie group structure of rotations and solves the problem of large-scale robust rotation averaging in a two-stage manner. -based optimization is first carried out to achieve an efficient, scalable, and robust solution. Based-on the -based initialization, Iterative Reweighted Least Squares (IRLS)-based optimization with the Geman-McClure (GM) loss function is performed to further improve the accuracy of the solution.
IRLS- [26] is the upgrade version of IRLS-GM, which also contains two steps. IRLS-GM and IRLS- share the same step one. For the second step of IRLS-, a fairly exhaustive evaluation of the performance of a variety of robust loss functions that are popular in the statistical literature is provided, which leads to a recommendation of the loss function that provides the best empirical performance.
MPLS [27] is a framework that could be served as alternative paradigm to IRLS for rotation averaging under high corruption and noise. It uses the theoretically guaranteed cycle-edge message passing algorithm for estimating the underlying corruption levels. These estimates are then used for learning better weights for the weighted least squares problem. The limitations of IRLS and the advantages of MPLS for solving the rotation averaging are also discussed and analyzed in the original paper of MPLS.
DESC [31] is a novel quadratic programming formulation for estimating the corruption levels in rotation averaging, which are used for solving this problem. Its objective function exploits the cycle consistency of the Lie group and has the following advantages: 1) It can tolerate corruption as high as the information-theoretic bound; 2) It does not require a good initialization for the estimates of group elements; 3) It has a simple interpretation; And 4) under some mild conditions its global minimum exactly recovers the corruption levels.
HARA [30] presents a hierarchical initialization scheme that constructs a spanning tree of a rotation graph by propagating most reliable constraints first and less reliable ones later. In HARA, the hierarchy of reliability based on the number of consistent triplet constraints, as well as their level of consistency, is established. That is, a constraint is considered to be more reliable if it is strongly supported by many other constraints and less reliable if it has weaker or fewer supports. In addition, the number of valid 2D-2D correspondences could also be optionally incorporated into the hierarchy.
NeuRoRA [28] is built to learn patterns from the data and predict/regress the model parameters from the noisy relative rotation measurements. NeuRoRA is a two-step approach: In the first step, a graph-based network, CleanNet, is utilized to clean the EG by removing outliers and rectifying noisy measurements; And in the second step, an initialization from the cleaned EG, instantiated from a shortest path tree, is then further fine-tuned using a separate graph-based network, FineNet, to produce the final rotation averaging results.
V-B Rotation Estimation Accuracy Benchmarking
For the methods of IRLS-GM, IRLS-, MPLS, DESC, and HARA, we locally run the official source codes (with default parameters) released by the authors on the new 1DSfM-derived rotation averaging benchmark to produce the absolute rotation estimation results of them. For the method of NeuRoRA, follow the authors’ instruction, the CleanNet and FineNet of NeuRoRA are both firstly pretrained on synthetic data and then finetuned and tested on the real-world data (with default parameters) in round-robin fashion (leave one out). And for the methods of existing IRA series (IRA, IRA++, IRAv3, and IRAv3+) and the proposed IRAv4, locally implemented source codes are employed for absolute rotation estimation. The absolute rotation estimation results of all the methods for comparison on the new benchmark are shown in Table II. The top-four results on each test data among all the comparative methods are highlighted, and in order to perform a comprehensive comparison, the rankings of all the comparative methods on each test data are averaged and shown in the last row of the table, and the top-four methods are highlighted as well. It could be seen from the table that: 1) Among all the rotation averaging methods for comparison, the proposed IRAv4 in this paper achieves overall the best performance in rotation estimation accuracy on the new 1DSfM-derived rotation averaging benchmark; 2) The deep learning-based rotation averaging method NeuRoRA is rather data/domain-sensitive and performs relatively poor in rotation estimation accuracy on the new benchmark, especially for the large-scale (PIC, ROF, and TFG) and high-noise (GDM, MDR, and USQ) test data; And 3) Compared with the results of IRA++ and IRAv3+, better performance is achieved by IRAv3 and IRAv4 in rotation estimation accuracy on the new benchmark, which demonstrates the effectiveness of the task-specific epipolar-geometry graph clustering and alignment reference construction proposed in IRAv3 and IRAv4, respectively.
In addition, in order to further demonstrate the effectiveness of the proposed IRAv4, comparative experiments in rotation estimation accuracy on the original 1DSfM dataset are also conducted, and the results are shown in Table III. It could be observed that compared with the state-of-the-art method on the dataset, RAGO [19], and the existing IRA series, IRAv4 is able to achieve overall the best performance as well. And the reasons for the inconsistency of different methods in rotation estimation accuracy on the two benchmarks (cf. Table II and Table III) mainly lie in the differences in ground-truth rotation source (different number of cameras involved in evaluation and different ways in obtaining their ground-truth rotations).
V-C Location Estimation Accuracy Benchmarking
Given the absolute rotation estimation results of all the rotation averaging methods for comparison on the new 1DSfM-derived benchmark (and the relative translations in the regenerated EG, which is mandatory for location estimation), we locally run the official source code of BATA (with default parameters) released by the authors to produce the absolute location estimation results of these methods. The absolute location estimation results of all the methods for comparison on the new benchmark are shown in Table IV. And similar to that in Table II, the top-four methods on each test data and in terms of overall rankings are highlighted. By observing the above two tables, similar conclusions, i.e. the three items discussed in the last sub-section, could be drawn. It could be further observed that though with moderately strong correlation, among different rotation averaging methods, the advantage in rotation estimation accuracy would not be strictly maintained during location estimation, e.g. the third and fourth-best methods in rotation estimation accuracy on the new benchmark are IRAv3 and IRLS-, while those in location estimation accuracy are IRA and IRLS-GM.
V-D Incorporation of the Global Bundle Adjustment
In order to demonstrate the performance of different rotation averaging methods coupled with some certain translation averaging method (BATA) in providing accurate enough camera pose initializations for the final BA, multi-view triangulation-based scene computation and global BA-based pose optimization are performed. Specifically, the pair-wise feature matches those pass through the RANSAC-based essential matrix estimation are linked to feature tracks with the union find algorithm implemented in OpenMVG111111https://github.com/openMVG/openMVG. Then, a RANSAC-based multi-view triangulation algorithm implemented in COLMAP is performed with the reprojection error and viewing angle thresholds, and the desirable level of estimation confidence setting to , , and to compute the initial values of the feature tracks’ spatial coordinates. During the multi-view triangulation, the intrinsic parameters provided by the original 1DSfM dataset, and the camera poses estimated by different rotation averaging methods and BATA are used to formulate camera projection matrices. Finally, given the initial values of camera parameters and scene points, global BA is performed to further improve the camera pose accuracy, which is conducted by leveraging Ceres Solver with Huber loss. The rotation and location estimation accuracy of the optimized camera poses of the top-2 methods in both rotation and location estimation accuracy before global BA (IRAv4 and MPLS) on the new benchmark are shown in the last two columns of Table II and Table IV, respectively. Note that the results of them are re-ranked independently in the tables to maintain comparative fairness. It could be observed from the two tables that with the global BA, the camera pose estimation accuracy of both IRAv4 and MPLS is largely improved, and IRAv4 achieves slightly better performance in location estimation accuracy than MPLS, which demonstrates that accurate enough camera pose initializations cloud be provided by both IRAv4 and MPLS coupling BATA, and the advantage of IRAv4 over MPLS.
Furthermore, to provide a more intuitive comparison in robustness of all the rotation averaging methods for comparison, the sparse scene reconstruction results after global BA of them on the test data of GDM of the new rotation averaging benchmark are shown in Fig. 2. As shown in the upper-left of this figure, there are two buildings with left-right symmetry located in the scene of GDM, which significantly increases the difficulty of camera pose estimation and the subsequent scene structure recovery. It could be observed from the figure that among all rotation averaging methods involved in benchmarking, only MPLS, DESC, and the IRA series succeed to recovery the correct scene structure of GDM, which demonstrate their robustness.
VI conclusion
In this paper, to benefit from the global reference for local-to-global rotation alignment more, an incremental parameter estimation-based task-specific connected dominating set extraction method is proposed, by which the accuracy and robustness of the Incremental Rotation Averaging (IRA) series are further advanced. Furthermore, a new 1DSfM-derived rotation averaging benchmark, together with new ground-truth data acquisition approach, and new evaluation tasks and metrics, based on the camera calibration results of COLMAP is presented, in order to give a more comprehensive and reasonable comparison among different rotation averaging methods. Based on the new benchmark, the proposed method in this paper, IRAv4, achieves overall the best performance in both rotation and downstream location estimation, which demonstrates its effectiveness in terms of accuracy and robustness when dealing with large-scale and high-noise rotation averaging problems.

### v4

@@ 4.2 Other key steps stemming from the IRAv3+
Absolute rotation global alignment and optimization: Given the alignment rotation of each cluster, the absolute rotations in the cluster’s local coordinate system are globally aligned to the reference’s coordinate system. After that, inlier edge set of the original EG is firstly obtained based on the aligned rotations and then used for providing constraints for globally optimizing them (cf. Eq. 11 and Eq. 12).
5 Experimental evaluation
In this section, the dataset and metric for evaluation are firstly described, and then the proposed IRAv4 method is evaluated in terms of absolute rotation estimation and alignment reference construction for its effectiveness demonstration.
5.1 Dataset and metric for evaluation
The evaluation dataset in this paper is the commonly-used 1DSfM dataset [45], which contains groups of test data stemmed from the Internet. The metadata of the dataset is listed in Table 1, where the absolute rotation ground truths of each test data are provided by the SfM results of Bundler [46]. The accuracy evaluation metric used in this paper is the same as that in the previous IRA series [29, 30, 31, 32]. Specifically, a best relative rotation between the absolute rotation estimations and ground truths is estimated and used to align these two rotation sets. And then the median alignment error in degree is used as the evaluation metric.
5.2 Comparison in absolute rotation estimation
Based on the introduced dataset and metric for evaluation, the performance of the proposed IRAv4 in absolute rotation estimation is validated in the manners of both synthetic and real-world experiments.
5.2.1 Comparison results on synthetic data
For the synthetic experiments, in order to create evaluation data that as close to the real-world one as possible, similar to CReTA [10], the synthetic data is created based on two test data of the 1DSfM dataset, MND and TOL. Specifically, the ground-truth relative rotations of all the epipolar edges are firstly obtained given the ground-truth absolute rotations. Then, following HARA [39] for each one of the relative rotation, a rotation disturbance whose rotation angle value being in accordance with a normal distribution is added. The standard deviation is set to either or . In addition, of the relative rotations are randomly selected and replaced by random rotations to simulate the outlier interruption. is set to , , , , , and . For comparison, 1 representative method for all 3 kinds of rotation averaging (robust loss-based, outlier filtering-based, and deep learning-based) methods is selected, including IRLS- [34], HARA [39], and NeuRoRA [40]. In addition, the primary method and latest one of the IRA series, i.e. IRA [29] and IRAv3+ [32], are also involved in the comparison. The synthetic experimental comparison results between our proposed IRAv4 and others are shown in Fig. 2.
It could be observed from the table that: On the one hand, under the same synthetic experiment configurations (noise level and outlier ratio), the rotation estimation accuracy on MND is higher than that of TOL on the whole; On the other hand, the accuracy drops when either the noise level or the outlier ratio increases. It should be further noted that the number of cameras is similar for MND and TOL ( vs. ), while the number of edges differs obviously ( vs. ). As a result, the hardness of the rotation averaging problem is highly related to noise level and graph density [47]. In addition, under all experiment configurations, our IRAv4 achieves best performance, which demonstrates its effectiveness in estimation accuracy and robustness to both estimation noise and measurement outliers.
5.2.2 Comparison results on real-world data
For the real-world comparative experiments, except for the previous IRA series, including IRA [29], IRA++ [30], IRAv3 [31], and IRAv3+ [32] which have been described in Sec. 3, IRAv4 is also compared with some other mainstream rotation averaging methods, including 3 robust loss-based methods, IRLS- [34], MPLS [35], DESC [36], 3 outlier filtering-based methods, OMSTs [37], HRRA [38], and HARA [39], and 3 deep learning-based methods, NeuRoRA [40], MSP [41], and RAGO [42]. The comparative results are illustrated in Table 2, and to deliver a comprehensive comparison, the rankings of the comparative methods on each test data are averaged and shown in the last row of the table. It could be observed that IRAv4 achieves overall best performance among all the comparative methods, which demonstrates its advantages in absolute rotation estimation.
In addition, to give a more comprehensive comparison on different rotation averaging methods, the absolute rotations obtained by them are fed into a representative translation averaging method for absolute camera location recovery. The methods for comparison are the same as those in the synthetic experiments, and the results are shown in Table 3, where the median alignment errors in camera locations wrt. the results of Bundler [46] are listed. It could be observed that by leveraging the absolute rotations provided by IRAv4, camera locations those with highest accuracy is achieved, especially for the most difficult test data, GDM, where symmetrical scene structure and numerous relative motion measurement outliers exist. Given the rotation and location estimates of GDM, multi-view triangulation [6] and global bundle adjustment [49] are performed for camera pose and scene structure optimization. And the sparse reconstruction results of GDM by IRLS-, HARA, NeuRoRA, and IRAv4-based global SfM are shown in Fig. 3. It could be observed that among the global SfM variants, only the one based on our IRAv4 could successfully recover the correct scene structure of GDM, which further validates its robustness to scenes with ambiguous structures and EG with high noise.
5.3 Comparison in alignment reference construction
As the most critical technique in IRAv4, the task-specific alignment reference construction method is compared with other two used in WCDS [33] and IRAv3+ [32], in terms of reference size and accuracy. The comparison results are shown in Table 3. It should be noted that for the method in IRAv3+, multiple CDSs are randomly extracted, by which different alignment references are constructed for different trails. Therefore, for the result of IRAv3+ in the table, trails reference construction by IRAv3+ are conducted and the mean value and standard deviation are listed. From the table we could see that: 1) On the one hand, compared with the naive method employed in WCDS [33], though reference with larger size is extracted by IRAv4, its accuracy is much higher; 2) On the other hand, compared with the method proposed in IRAv3+, though with similar values in both and , the reference constructed by IRAv4 is with smaller size and higher accuracy on the whole. As a result, the effectiveness of the proposed reference construction method is validated.
6 Conclusion
In this paper, to benefit from the global alignment reference for local-to-global rotation alignment more, an incremental parameter estimation-based task-specific connected dominating set extraction method is proposed, by which the accuracy and robustness of the Incremental Rotation Averaging (IRA) series are further advanced. The proposed method in this paper, IRAv4, is comprehensively evaluated on the 1DSfM dataset and achieves overall the best performance in both absolute rotation estimation and alignment reference construction, which demonstrates its effectiveness when dealing with large-scale and high-noise rotation averaging problems.
CRediT authorship contribution statement
Xiang Gao: Conceptualization, Formal analysis, Methodology, Software, Validation, Writing – original draft. Hainan Cui: Data curation, Investigation. Yangdong Liu: Project administration, Visualization. Shuhan Shen: Funding acquisition, Resources, Supervision, Writing – review & editing.
Acknowledgements
This work was supported by the National Science Foundation of China under Grant 62373349, Grant U22B2055, and Grant 62402495, and the Key R&D Project in Henan Province under Grant 231111210300.
Declaration of competing interest
The authors have no competing interests to declare that are relevant to the content of this article.
Data availability
Data will be made available on request.